Random samples of sizes n1 = 32 and n2 = 40 are to be drawn from two independent populations.

μ1 = 12.3      μ2 = 9.8

σ1 = 2.9         σ2 = 2.4

  1. P(Xbar1 - Xbar 2 < 2)
  2. P(S12/ S22 > 2)
Answer #1

a.

P(\bar{X_1}-\bar{X_2}<2)=P(\frac{(\bar{X_1}-\bar{X_2})-(\mu_1-\mu_2)}{\sqrt{\frac{\sigma_1^2}{n_1}+\frac{\sigma_2^2}{n_2}}}<\frac{2-(12.3-9.8)}{\sqrt{\frac{2.9^2}{32}+\frac{2.4^2}{40}}})

=P(Z<\frac{2-2.5}{0.6378})=P(Z<-0.78)=\Phi(-0.78)=1-\Phi(0.78) , Z\to N(\mu=0,\sigma=1)

=1-0.7823=0.2177 ; From standard normal distribution table

b.

P(\frac{s_1^2}{s_2^2}>2)=P(\frac{\frac{(n_1-1)s_1^2}{\sigma_1^2}}{\frac{(n_2-1)s_2^2}{\sigma_2^2}}>2*\frac{\frac{n_1-1}{\sigma_1^2}}{\frac{(n_2-1)}{\sigma_2^2}})

=P(\frac{\chi^2_{n_1-1}}{\chi^2_{n_2-1}}>2*\frac{\frac{32-1}{2.9^2}}{\frac{(40-1)}{2.4^2}})

=P(\frac{\frac{\chi^2_{n_1-1}}{n_1-1}}{\frac{\chi^2_{n_2-1}}{n_2-1}}>2*\frac{\frac{32-1}{2.9^2(32-1)}}{\frac{(40-1)}{2.4^2(40-1)}})

=P(F_{n_1-1,n_2-1}>2*\frac{2.4^2}{2.9^2})

=P(F_{32-1,40-1}>1.3698)

=P(F_{31,39}>1.3698)

=0.1750 ; The Excel function is , =FDIST(1.3698,31,39)